OpenStudy (anonymous):

Will give medal What is 28% of 75? A. 37 B. 21 C. 28 D. 47

OpenStudy (narissa):

21

OpenStudy (anonymous):

do 28% times 75 and you will get B

OpenStudy (solomonzelman):

you are taking 28% of 75. this means that you are taking 28/100 th of it. (To find for example 3/4th of 75 you would multiply 75 by 3/4)

OpenStudy (ilovecake):

b

OpenStudy (anonymous):

oooh ok thanks guys

OpenStudy (solomonzelman):

So, any time you want to find \(\large\color{black}{ \displaystyle \text{A%} }\), of a certain number \(\large\color{black}{ \displaystyle {\rm D} }\), you would do the following: \(\large\color{black}{ \displaystyle {\rm D} \times\frac{\text{A} }{100} }\)

OpenStudy (anonymous):

oh ok

OpenStudy (solomonzelman):

usually: (A can be larger than 100, but has to be a positive number) (D can be any positive number)